OpenStudy (anonymous):

Tell me if I am right Solve for x 2x^2=-13x-20 answer: x=-4,-5/2

OpenStudy (anonymous):

\[2x^2=-13x-20\]\[2x^2+13x+20=0\]\[(x+4)(2x+5)=0\]\[x=-4 \cup x=-5/2\]I agree with your answers :})